What is the angle of elevation of the sun when a 35-ft mast casts a 20-ft shadow?
Found 2 solutions by stanbon, Nate:
Answer by stanbon(48558)

What is the angle of elevation of the sun when a 35-ft mast casts a 20-ft shadow?
------------
Draw the picture.
You have a right triangle with sides 35 and 20 ft.
The 35 is opposite the angle of elevation; the 20
is adjacent to the angle of elevation.
arctan(35/20)= 60.25 degrees or 1.05 radians.
Cheers,
Stan H.

Answer by Nate(3500)

Think of this:
The shadow makes a leg of a triangle. The height of the mast makes the other leg.
Adj. leg ~> 20
Opp. leg ~> 35
Tan( angle of elevation ) = opp. / adj. = 35/20 = 7/4
angle of elevation = cot( 7/4 ) or about 60.255118....
Angle of elevation is roughly 60 degrees.